[Brahms-l] Shift report 20040104 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 04 2004 - 23:55:12 EST
------------------------------------------------------------
Date: Sun Jan  4 23:55:11 2004, Shift: 16:00-24:00
Supervisor: sjs, with: 
Log book#15 pages 23 - 23

Shift Summary:
------------------------------------------------------------
Good beam for latter part of shift.  
With current stores we are
finding our background getting worst when they collimate 
the beam, but the uncollimated ZDC/TOFW ratio looks very good to start with.
Started shift with pedestal runs 9360 and 9361.  Significant
data runs 9363 (FS 8deg, 1/8B; MRS 90 deg, A150){248K trig6),
9364,9365,9367 (FS 1/8A){350K trig6}.  Started run 9368. No new problems detected.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 4 23:55:30 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 04 2004 - 23:55:48 EST